OpenStudy (matt6288):

The problem is Sec^2-Tan^4=(1+Sin^2)/Cos^2 I have to determine if its true or not

OpenStudy (matt6288):

I don't know what to do to each side? 1/Cos^2x - Sin^4x/Cos^4x

OpenStudy (freckles):

just to make sure I know what you said I'm going to type it pretty latex math stuff \[\sec^2(x)-\tan^4(x)=1+\frac{\sin^2(x)}{\cos^2(x)} \]

OpenStudy (lυἶცἶ0210):

It is that right? ^ Not (1+s^2)/c^2?

OpenStudy (matt6288):

=(1+sin^2x)/cos^2x

OpenStudy (freckles):

try to plug in a number and see if both sides are the same for that number if they are it could be an identity and we can try to see if we can prove both sides are the same try enter in for x pi/4

OpenStudy (matt6288):

I tried multiplying the left side by cos^2

OpenStudy (matt6288):

than i got sin^4/cos^2

OpenStudy (matt6288):

\[ \frac{ 1}{ \cos^2\theta }-\frac{ \sin^4\theta }{\cos^4\theta }\]

OpenStudy (freckles):

I guess you tried pluggin in pi/4 ?